OpenStudy (anonymous):

-x^5-x^3+6 how do you factor this?

OpenStudy (anonymous):

i think it might be easier to factor \[-(x^5+x^3-6)\]

OpenStudy (anonymous):

this is what I got -x(x^4+x^2+6). is it right?

OpenStudy (anonymous):

or maybe not not clear how to factor this at all

OpenStudy (anonymous):

no, not unless you had a typo in the problem

OpenStudy (anonymous):

-x^5-x^3+6x oops yea sorry it had 6x

OpenStudy (anonymous):

could it have been \[-x^5-x^3+6x\]?

OpenStudy (anonymous):

yes, I messed up

OpenStudy (anonymous):

ok i see your answer was close, but not quite should have been \[ -x(x^4+x^2-6)\]

OpenStudy (anonymous):

then you can continue to factor, since \(x^4+x^2-6\) factors

OpenStudy (anonymous):

I'm didn't know you could keep factoring..? I'm trying to find the x intercept afterwards, could you help me walk through this one?

OpenStudy (anonymous):

sure if it was \(x^2+x-6\) you could probably factor it easy right?

OpenStudy (anonymous):

right that would have been (x-2)(x+3)

OpenStudy (anonymous):

right, and this is exactly the same, except instead of \(x^2\) and \(x\) you have \(x^4\) and \(x^2\) noting that \(x^4=(x^2)^2\)

OpenStudy (anonymous):

so it factors in exactly the same way, but with \(x^2\) instead of \(x\) i.e. \[x^4+x^2-6=(x^2-2)(x^2+3)\]

OpenStudy (anonymous):

final answer therefore looks like \[-x(x^2-2)(x^2+3)\]

OpenStudy (anonymous):

oh, so you would ignore the -x on the outside first and factor what's in the parenthesis. And to find the x-int. you would set y=o right? f(x)=-x(x^2-2)(x^2+3) 0=-x(x^2-2)(x^2+3) x=2 x=-3 ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[0=-x(x^2-2)(x^2+3)\] is right

OpenStudy (anonymous):

then \(-x=0\iff x=0\) is also right, but \[x^2-2=0\] is not \(x=2\)

OpenStudy (anonymous):

\[x^2-2=0\\ x^2=2\\ x=\pm\sqrt2\] is more like it

OpenStudy (anonymous):

\[x ^{2}+3=x=\pm \sqrt{-3}=\pm i \sqrt{3}?\]

OpenStudy (anonymous):

how would you graph these as a sq root?